Ali has 18new magazines to read.
Let M be the number of magazines he would have left to read after reading R of them.

Write an equation relating M to R. Then use this equation to find the number of magazines he would have left to read after reading 7 of them.
1
Expert's answer
2020-03-02T15:05:13-0500

M = Number of unread magazines

R= Magazines already read

we need to find number of unread magazines (M) after reading 7 magazines (i.e. R = 7)


Unread magazines + magazines already read = Total number of new magazines

M + R = 18

Subtracting R from both sides

M + R - R = 18 - R

M = 18 - R ..............required equation.

substituting R= 7

M = 18 - 7

M = 9

So there will be 9 magazines left to read.